# A regularized continuum model for traveling waves and dispersive shocks of the granular chain

### v1

@@ 9 Conclusions and future challenges
Acknowledgments
The authors thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “Emergent phenomena in nonlinear dispersive waves” held in Newcastle in July–August, 2024. We also thank Gennady El for valuable conversations. This material is based upon work supported by the U.S. National Science Foundation under the awards DMS-2107945 (C.C.), PHY-2110030, PHY-2408988 and DMS-2204702 (PGK).
Appendix A
In this appendix A, we display all the detailed steps of the computation of periodic wave solutions of the continuum PDE model (2.12).
A.1 Periodic solutions
We showcase the detailed calculation procedures to obtain the periodic wave solutions of Eqs. (4.3) and (4.5).
A.1.1 $p=3/2$
@@ A.2 $p=2$
Appendix B
In this appendix, we present some details about the derivation of the Whitham modulation equations for the case of generic values of $p$ for the continuum model (2.12).
With the co-traveling frame transformation (i.e., $r\left(X,T\right)=R(\theta)$, with $\theta=(KX-\Omega T)/\epsilon$), we obtain that, after the substitution, the function $R(\theta)$ satisfies the differential equation
$\left(R_{\theta}\right)^{2}=G\left(R\right),$ (B.1)
where
$G\left(R\right)=\frac{12}{K^{2}}\left(-\frac{2}{c_{p}^{2}\left(p+1\right)}R^{p+1}+R^{2}-2BR-C\right),$ (B.2)
$B,C$ being constants of integration and $c_{p}=\Omega/K$. For generic values of $p$, the solution to the ODE (B.1) is not analytically tractable. Importantly, however, we expect that the solutions will still form a 3-parameter family. This is because, even though the degree of the right-hand side of Eq. (B.1) depends on $p$, it always contains the same number of parameters, namely $B,C,K,\Omega$. By demanding that the periodic wave in terms of $\theta$ is $2\pi$ periodic, this reduces the total number of parameters to three. While we will obtain one more integration constant, since Eq. (B.1) is a first-order differential equation, this will be the phase shift which we can set to 00 without loss of generality. Therefore, the number of independent parameters in the solutions is independent of $p$.
To derive the Whitham modulation equations, we notice first that the Lagrangian density for the continuum model at the level of the displacement (5.3) is now given as follows:
$\mathbb{L}=\frac{1}{2}\left(u_{T}\right)^{2}+\frac{\epsilon^{2}}{24}u_{XX}u_{TT}-\frac{1}{p+1}\left(u_{X}\right)^{p+1},$ (B.3)
where $u_{X}=r$. With these changes, the derivation of the modulation equations proceeds as in section 6. Therefore, we omit the details for brevity.

### v2

@@ 9 Conclusions and future challenges
It is relevant to once again point out that the applicability of the DSW fitting method requires the corresponding modulation equations be genuinely nonlinear and strictly hyperbolic, e.g., see [22, 48, 45, 46], and it is not known at present whether the modulation equations derived in section 6 enjoy these properties. When the modulation equations are not genuinely nonlinear and strictly hyperbolic, the trailing edge speeds can in some cases display a non-monotonic dependence on the jump height, e.g., see [48] for the Serre-Green-Naghdi equations. In the present case, however, the numerically computed trailing edge speeds for both the discrete and continuum model display a monotonic dependence on the size of the jump, even though the genuine nonlinearity and strict hyperbolicity of the modulation equations are presently unknown and constitute an interesting open problem for future studies.
@@
Appendix A: Further details on the derivation of the periodic solutions
In this appendix we display the detailed steps of the computation of the periodic wave solutions of the continuum PDE model (2.12), specifically, Eqs. (4.3) and (4.5).
A.1 $p=3/2$